(1) Omar spent the first part of the night working on the guiding problem at rotator angles other than 270deg. Guiding at rotator angles of 0 and 180 deg was tested. Guiding was much better after Omar's work.
(2) It would help if the TCS console for the operators had a display that would report the parallactic angle for a given object at a given time. We used the parallactic angle display from the COSMOS GUI. If there is a parallactic angle display and calculator in the Blanco TCS, we could not find it.
(3) We has two or three errors in the COSMOS DHS system. The message in the DHS vncviewer reported "DHS SMCMgr cosmosdhs waiting for data" and the error in the xterm window from which we launced the observing software stated" "dcaRecv read error 104. Connection reset by peer DITS exited with signal SIGPIPE"
(4) I do not know if this is a problem or not. The vncwindow that we use to launch and run the observing software on observer2 was very sluggish today. We also noticed that observer1 was covered with windows for DECam. Do all of these DECam windows need to be running and/or displayed when DECam is not in use?
(5) COSMOS is not present as an option in the drop down menu of this night report form.
Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
---|---|---|---|---|---|---|
2014-07-11 | 11.25 | 0.0 | 0.0 | 0.0 | 0.0 | 11.25 |
Overall we had a very good night and a productive engineering block. We used the old RCSpec calibration system that was installed on 2014-07-10 for our calibrations. During the afternoon, openings in the Cass rotator were covered with cardboard and tape.
We took spectra of a standard star (EG274) at the parallactic angle and parallactic angle + 90deg with the ADC in. This preliminary test suggests that the ADC is working as expected. We also observed some SNe from the ASASSN group.
This engineering run has/had three major goals to (1) test the ADC to make sure that it works, (2) be able to rotate CO SMOS at Cass so that we can align MOS masks or match the long slits to the parallactic angle, and (3) have a working calibration lamp system. The ADC is working, the rotator has the ability to be moved in course adjustment mode (accuracy of a few tenths of a degree). The fine adjusment encoders for the rotator will be completed before the Oct COSMOS run. The calibration unit allows us to take arc spectra for wavelength calibrations. We need to place a quartz lamp into the calibration unit so that we can obtain quartz flats. We have noticed that the arc lamps do not appear to be uniformly illuminating the slit. This may be something to work on as well.
Many thanks to everybody who has helped us before, during, and after this run.